OpenStudy (anonymous):

If sin θ = 12 over 37, use the Pythagorean Identity to find cos θ.

Nnesha (nnesha):

Identity is \[\huge\rm sin^2 \theta +\cos^2 \theta =1\] substitute sin for 12/37 solve for cos

Nnesha (nnesha):